A simple pendulum of length 60.0 cm takes to make 12 complete back and forth motion.


(A) find it's period. (B) what will be it's period when it's Length is increased to 240.0 cm

1
Expert's answer
2021-11-29T11:47:00-0500

(A) "T=2\pi\sqrt{l/g}=2\cdot3.14\cdot\sqrt{0.6/9.8}=1.55\ (m)"


(B) "T=2\pi\sqrt{l/g}=2\cdot3.14\cdot\sqrt{2.4/9.8}=3.11\ (m)"